OpenStudy (anonymous):

A ball has a diameter of 8 inches. What is its surface area? (Use 3.14 for π.)

OpenStudy (lgbasallote):

\[SA = 4\pi r^2\] \[radius = \frac{diameter}{2} = \frac{8}{2} = 4 in\] therefore \[SA = 4(3.14)(4)^2\] got it?

OpenStudy (anonymous):

i believe the answer would be 1607.60 if you would like to check it here the website http://www.icoachmath.com/solvedexample/sampleworksheet.aspx?process=/__cstlqvxbefxaxbgefdxkjhfbxedbaexd&.html Guudluck!(:

OpenStudy (anonymous):

my choices are -200.96 in.2 803.88 in.2 401.92 in.2 100.48 in.2

OpenStudy (lgbasallote):

@miadeex calculate the one i gave you ;) it'll give you one of those

OpenStudy (anonymous):

lol how do i do that

OpenStudy (anonymous):

SA= 4*3.14*(4*4)

OpenStudy (lgbasallote):

multiply \[4 \times 3.14 \times 4^2\]

OpenStudy (lgbasallote):

you can use google, calculator, wolframalpha..or by hand..your choice ;)

OpenStudy (anonymous):

since the answers choices are not close to each other you can estimate the answer by simplifying the problem. 4*3*4*4

OpenStudy (anonymous):

Your answer would be 803.84 trust me(;

OpenStudy (anonymous):

The closest answer to that is what you get

OpenStudy (lgbasallote):

@JoannaiA i believe you did 8^2..not 4^2

OpenStudy (anonymous):

and i believe yu r most certanly correct i did it wrong would it be choice A then? sorry tryed helping

OpenStudy (lgbasallote):

@miadeex what do you think would be the answer?

OpenStudy (lgbasallote):

and it's alright joanna :) let's try to see what mia got first then we'll discuss the answer

OpenStudy (anonymous):

as said by @JoannaiA you can estimate it by 4 x 3 x 4 x 4 = around 200 so ur answer should be 200.96

OpenStudy (anonymous):

thanks guys you all helped a ton it was 200.96

OpenStudy (lgbasallote):

yep it is